# Charm physics with overlap fermions on 2+1-flavor domain wall fermion configurations

### v1

@@ 4 Summary
In this work, we have computed decay constants and their ratios for $D$, $D^{*}$, $D_{s}$, $D_{s}^{*}$, $\eta_{c}$ and $J/\psi$, using $2+1$-flavor configurations. We have also calculated the decay constants with tensor currents of vector mesons and provided estimates of discretization errors. Our final results for the decay constants adding up all errors are given in Tab. 12 (in units of MeV), of which the ratios are given in Tabs. 10 and 11 for charmed mesons, and in Eq. (23) for charmonia. The ratios in Tabs. 10 and 11 reflect the magnitudes of heavy quark symmetry breaking and SU(3) flavor symmetry breaking, respectively. The precision of $f_{D_{s}^{*}}^{T}/f_{D_{s}^{*}}$ is improved compared with that from our previous work [26]. We compare $f_{D^{*}_{(s)}}$ from this work and other lattice QCD calculations in Tab. 13.
Table 12: Decay constants in units of MeV. The ones with tensor currents are renormalized in the ${\overline{\rm MS}}$ scheme at the scale of 2 GeV.

### v3

@@ 4 Summary
In this work, we have computed decay constants and their ratios for $D$, $D^{*}$, $D_{s}$, $D_{s}^{*}$, $\eta_{c}$ and $J/\psi$, using $2+1$-flavor configurations. We have also calculated the decay constants with tensor currents of vector mesons and provided estimates of discretization errors. Our final results for the decay constants adding up all errors are given in Tab. 13 (in units of MeV), of which the ratios are given in Tabs. 11 and 12 for charmed mesons, and in Eq. (23) for charmonia. The ratios in Tabs. 11 and 12 reflect the magnitudes of heavy quark symmetry breaking and SU(3) flavor symmetry breaking, respectively. The precision of $f_{D_{s}^{*}}^{T}/f_{D_{s}^{*}}$ is improved compared with that from our previous work [34]. We compare $f_{D^{*}_{(s)}}$ from this work and other lattice QCD calculations in Tab. 14.
Table 13: Decay constants in units of MeV, among which the $f_{\eta_{c}}$ and $f_{J/\psi}^{T}$ require $f_{J/\psi}$ as an additional input. The ones with tensor currents are renormalized in the ${\overline{\rm MS}}$ scheme at the scale of 2 GeV.